### v2

# MAESTRO: Multi-Agent Environment Shaping through Task and Reward Optimization

Abstract
Cooperative multi-agent reinforcement learning (MARL) is constrained by two core design bottlenecks: dense reward engineering and curriculum design. Hand-crafted rewards often misalign with global objectives, while static curricula fail to prevent agents from getting trapped in local optima. Large language models (LLMs) have been used to address some of these issues, but most work adopts an LLM-as-Agent paradigm that keeps the model in the control loop, incurring prohibitive latency and token costs for real-time systems.
We propose MAESTRO (Multi-Agent Environment Shaping through Task and Reward Optimization), which shifts the LLM’s role from online controller to offline training architect. MAESTRO treats training as a dual-loop optimization problem: an LLM shapes the learning environment, while a standard MARL backbone optimizes the policy. The LLM instantiates two generative components: (i) a semantic curriculum generator that synthesizes diverse, performance-conditioned traffic scenarios, and (ii) an automated reward and policy synthesizer that fills executable Python templates to produce reward shaping signals and prior policy logits. Prior policies are used only as a regularizer during actor updates (via MSE penalties that decay over training); the deployed controller is a standalone MADDPG policy with no LLM calls. Reward shaping is additively combined with environment rewards, with a difficulty-dependent weight.
[…]
Code is available at https://github.com/BrianWu1010/MAESTRO.git.
1 Introduction
Cooperative multi-agent reinforcement learning (MARL) is a natural framework for decentralized control problems such as warehouse logistics, sensor networks, and urban traffic signal control (TSC). In principle, centralized training with decentralized execution allows agents to coordinate via learned policies while acting on local observations. In practice, scaling MARL to realistic, non-stationary environments remains difficult due to two persistent design challenges: curriculum design and reward engineering.
Curriculum learning (CL) addresses exploration by exposing agents to gradually harder tasks. However, effective curricula require domain knowledge to ensure that tasks are both diverse and aligned with the target deployment regime [narvekar2020curriculum]. Reward design is equally critical: dense rewards may mis-specify long-term objectives [zang2020metalight], while sparse global rewards provide weak learning signals. Crafting rewards that generalize across dynamic regimes is labor-intensive and brittle.
LLMs offer rich semantic priors that could automate both curricula and rewards. Yet most existing integrations follow an LLM-as-Agent paradigm: the LLM directly selects actions or high-level decisions in the control loop [lai2023llm, da2023llm, zhang2023trafficgpt]. This can yield strong zero-shot behavior but incurs high latency and token costs, limiting its applicability to real-time, safety-critical systems. It also underuses the LLM’s strength as a high-level designer.
Our approach: LLM-as-Architect.
We propose MAESTRO (Multi-Agent Environment Shaping through Task and Reward Optimization), which repositions the LLM as an offline architect of the training process. Rather than emitting actions, the LLM designs the training environment: it generates curriculum contexts and fills parameterized templates for reward shaping and prior policy logits. A standard MADDPG backbone then optimizes a policy under this shaped environment and reward.
MAESTRO’s training loop has two coupled components:
  1. 1.
    A semantic curriculum generator that maps a scalar difficulty level and performance statistics into natural-language traffic descriptions, which are compiled into CityFlow configurations.
  2. 2.
    An automated reward and policy synthesizer that outputs parameters for pre-validated Python templates. These templates implement an auxiliary reward function and a prior policy that regularizes the actor via a time-decaying MSE penalty, without intervening in action selection.
The resulting policy remains lightweight and low-latency at deployment, while the LLM’s reasoning is fully exploited during training.
Figure 1: MAESTRO workflow. The LLM (Architect) generates curriculum contexts and parameterizes executable templates for reward shaping and prior policy logits. MARL agents (Learners) train under this shaped environment using MADDPG. Performance feedback drives curriculum updates and, in A7, periodic regeneration of reward and policy templates. LLM calls occur only during training, not deployment.
Contributions.
We validate MAESTRO on urban TSC, a standard cooperative MARL benchmark [wei2019colight, zhang2019cityflow], and make three contributions:
  • Generative environment shaping. We formalize MAESTRO as a contextual MARL framework in which an LLM automates both task generation and reward synthesis while remaining entirely offline at deployment.
  • Controlled ablation of LLM components. We compare three configurations: A2 (adaptive LLM curriculum with template rewards), A8 (alternative LLM curriculum with template rewards), and A7 (full MAESTRO with adaptive curriculum and LLM-shaped rewards), isolating the effects of curriculum design and reward shaping.
  • Stability and risk-adjusted performance. Across four seeds, A7 achieves markedly higher stability than both A2 and A8 while still improving mean return. In domain metrics, MAESTRO configurations reduce delay by 20–28% and increase throughput by 6–9% relative to the baseline.
2 Related Work
2.1 MARL for Decentralized Control
MARL is widely used for cooperative control with a shared global objective and local observations. Centralized training with decentralized execution (CTDE) [lowe2017maddpg, rashid2018qmix] mitigates non-stationarity during learning by training joint critics while deploying decentralized actors. Even under CTDE, state-of-the-art methods can be sensitive to configuration and hyperparameters [papoudakis2021benchmarking], motivating additional structural priors.
Urban TSC is a canonical MARL benchmark. CoLight [wei2019colight] and MPLight [chen2020mplight] use graph attention and pressure-based features to model inter-agent interactions and achieve strong performance on real networks. However, they assume fixed environment specifications and manually designed rewards. Reward shaping theory [ng1999policy] warns that naive shaping can induce suboptimal policies, and in practice TSC rewards remain heuristic. MAESTRO retains a CTDE backbone (MADDPG) but allows both curricula and rewards to evolve via LLM-generated signals.
2.2 Automatic Curriculum and Environment Generation
Curriculum learning sequences tasks by difficulty to accelerate training [bengio2009curriculum]. Automatic curriculum learning (ACL) extends this idea by adapting task difficulty based on learning progress [portelas2020automatic]. Adversarial and co-evolutionary approaches such as PAIRED [dennis2020emergent] and POET [wang2019paired] construct challenging environments that drive robust skill acquisition.
In cooperative MARL, difficulty-aware curricula (e.g., cMALC-D [kim2023cmalcd]) adjust task parameters based on performance. In TSC, curricula typically perturb numeric parameters such as arrival rates or demand scales [zang2020metalight], which improves learning but captures only a narrow slice of real-world structure (e.g., complex rush-hour patterns or localized bottlenecks).
MAESTRO extends ACL in two ways. First, it leverages an LLM to generate semantically rich traffic descriptions, which are compiled into CityFlow configs, expanding the curriculum space beyond low-dimensional numeric perturbations. Second, reward shaping and curriculum progression are coupled: as difficulty changes, the LLM simultaneously revises environment descriptions and reward logic, aligning training signals with the evolving task distribution.
2.3 LLMs for Reward and Training-Process Generation
Two main paradigms integrate LLMs with RL.
LLM-as-Agent (online control).
Here, the LLM selects actions or high-level plans at inference time. TrafficGPT [zhang2023trafficgpt] and LLM-TSCS [lai2023llm] apply this to traffic management; LA-Light [lai2024lalight], Voyager [wang2023voyager], and Code-as-Policies [liang2023code] do so in robotics and embodied environments. These systems exploit LLM priors for zero-shot generalization but inherit high latency, token costs, and limited action frequency, making them problematic for real-time multi-intersection TSC.
LLM-as-Teacher (offline guidance).
An alternative is to use LLMs to shape training rather than deployment. Eureka [ma2023eureka] prompts an LLM to generate and refine reward functions for robotic tasks, improving policies without using the LLM at test time. In traffic control, Da Lio et al. [da2023llm] inject semantic priors via language, and LAMARL [lin2024lamarl] uses language-based heuristics to guide MARL with curriculum hints and shaping.
Within this teacher regime, an important distinction is whether the LLM remains in the deployed controller. MAESTRO belongs to the class where the runtime agent is purely an RL policy: the LLM is used only during training to generate curricula, reward functions, and policy priors. Compared to prior work, MAESTRO (i) applies template-based LLM reward synthesis to a cooperative 16-agent TSC setting, and (ii) tightly couples LLM-shaped rewards with an adaptive, performance-driven curriculum. We show that this joint environment–reward generation is key to the stability and risk-adjusted gains observed.
3 Preliminaries
3.1 Contextual MARL Formulation
We model multi-intersection TSC as a contextual Decentralized Partially Observable Markov Decision Process (Dec-POMDP), augmented with curriculum contexts generated by the LLM Architect:
$\mathcal{M}_{C}=\langle\mathcal{C},\mathcal{I},\mathcal{S},\{A_{i}\}_{i\in\mathcal{I}},\{O_{i}\}_{i\in\mathcal{I}},\mathcal{T}_{C},\mathcal{R}_{C},\gamma\rangle,$
where:
  • $\mathcal{C}$ is the curriculum context space (LLM-generated traffic patterns).
  • $\mathcal{I}=\{1,\dots,N\}$ indexes intersections.
  • $\mathcal{S}$ is the global state; $O_{i}$ and $A_{i}$ are the observation and action spaces for agent $i$.
  • $\mathcal{T}_{C}$ is the context-dependent transition function.
  • $\mathcal{R}_{C}$ is the context-dependent reward function; $\gamma$ is the discount factor.
At each training stage, a context $C\in\mathcal{C}$ is sampled, and the LLM synthesizes $\mathcal{R}_{C}$. Agents then follow a joint policy $\pi_{C}$ to generate a trajectory $\tau$. The objective is to maximize expected return over the curriculum distribution:
$\max_{\pi}\;\mathbb{E}_{C\sim p(\mathcal{C}),\tau\sim\mathcal{T}_{C}(\cdot|\pi_{C})}\left[\sum_{t=0}^{T}\gamma^{t}\mathcal{R}_{C}(s_{t},a_{t})\right].$ (1)
We adopt MADDPG [lowe2017maddpg] as the MARL learner under CTDE, ensuring stable cooperative learning and fast decentralized execution.
3.2 Environment and Agent Specification
Simulation setup.
We use CityFlow [zhang2019cityflow] on the real-world Hangzhou (HZ) network with 16 signalized intersections. Each episode spans 360 steps (1-second simulation interval). Context $C$ controls route distributions, arrival rates, and vehicle parameters; ranges are detailed in the appendix.
Observations and actions.
Each agent $i$ observes a vector $o_{i}^{t}$ including local queue lengths, inflow/outflow, active phase, elapsed phase duration, occupancy, and short-term history, following standard MARL-TSC practice. Actions select one of four feasible signal phases per intersection.
Traffic metrics.
We track mean travel time, delay, queue length, and throughput. These metrics both guide curriculum difficulty updates and form components of the composite reward $\mathcal{R}_{C}$ used for training.
4 Methodology
Figure 1 illustrates MAESTRO’s dual-loop training. The outer loop updates curriculum difficulty and (in A7) triggers LLM regeneration of reward and policy templates; the inner loop runs standard MADDPG updates under the current shaped reward.
Figure 2: Conceptual foundation of MAESTRO. The LLM (Architect) generates task-specific components—curriculum contexts, reward parameters, and prior policy parameters—that are injected into the MARL training loop. The deployed controller is an unmodified MADDPG policy; LLM inference is restricted to training.
4.1 Configurations
We evaluate three configurations that share the same network architecture, hyperparameters, and simulator, differing only in their use of the LLM:
Config Curriculum Mode Reward Source LLM Reward Shaping?
A2 llm_adaptive Template-based (environment) No
A8 llm Template-based (environment) No
A7 llm_adaptive Environment + LLM-shaped Yes
All settings use LLM-generated curricula; only A7 adds LLM-shaped reward and policy priors.
Curriculum modes.
  • llm_adaptive (A2, A7): Structured prompts with explicit difficulty mapping and conservative regeneration triggers; parameters are clipped to bounded ranges.
  • llm (A8): Less constrained prompts allowing more diverse contexts, increasing expressiveness but also variability.
Both modes update a scalar difficulty $d\in[0.3,1.0]$, but llm tends to produce more volatile curriculum trajectories.
4.2 Adaptive Curriculum
The curriculum maintains a scalar difficulty $d\in[0.3,1.0]$, updated from a rolling window of five episodes using cumulative return. Let $\bar{R}_{w}$ denote the windowed mean and $\tau=165.0$ a fixed threshold. The update rule is:
[…]
The asymmetric patience (2 for increases, 1 for decreases) favors stability by requiring sustained success before progression while quickly reacting to deterioration. Difficulty bounds avoid trivial or saturated regimes.
Under llm_adaptive, GPT-4o-mini receives the current $d$ and summary statistics and outputs a natural-language traffic context; this is compiled into CityFlow config files. Under llm, prompts are less constrained but share the same difficulty updates. Pilot experiments under a fixed-time controller confirm that increasing $d$ monotonically increases difficulty (queue length, delay).
The threshold $\tau=165.0$ corresponds to the mean return of a competent fixed-time baseline over 200 episodes and is stable (variance <<3%). Multipliers $1.05$ and $0.95$ create hysteresis around this anchor.
Difficulty is mapped to CityFlow parameters via linear interpolation; parameters that increase difficulty (e.g., maximum speed, acceleration) grow with $d$, while others (e.g., minimum gap, headway) shrink. Each parameter is further perturbed by $\pm 20\%$ multiplicative noise to diversify scenarios at fixed difficulty.
4.3 Training Algorithm
Algorithm 1 summarizes MAESTRO. The outer loop samples curriculum tasks and updates difficulty; the inner loop performs MADDPG updates using either environment-only rewards (A2/A8) or combined environment + LLM shaping (A7).
Algorithm 1 MAESTRO: Dual-Loop Training with Adaptive Curriculum and LLM Guidance
1:Initialize difficulty $d\leftarrow 0.3$, window $W\leftarrow\emptyset$, policy $\pi_{\theta}$, replay buffer $\mathcal{D}$
2:Initialize LLM reward $r_{\text{LLM}}\leftarrow\text{None}$, prior policy $\pi_{\text{LLM}}\leftarrow\text{None}$
3:Set threshold $\tau=165.0$, patience $(p^{+},p^{-})=(2,1)$, regularization schedule $\alpha(t)$
4:while $t_{\text{env}}<T_{\text{max}}$ do $\triangleright$ Outer loop
5:  task $\leftarrow\text{GenerateTask}(d)$ $\triangleright$ LLM-generated context
6:  Reset environment with task; $\tau_{\text{episode}}\leftarrow\emptyset$
7:  for $t=1$ to $T_{\text{episode}}$ do
8:    $a_{t}\sim\pi_{\theta}(s_{t})$ with exploration noise
9:    $r_{\text{env},t},s_{t+1}\leftarrow\text{Environment.step}(a_{t})$
10:    if $r_{\text{LLM}}\neq\text{None}$ then $\triangleright$ A7 only
11:     $w\leftarrow 0.1+0.4\cdot d$
12:     $r_{t}\leftarrow r_{\text{env},t}+w\cdot r_{\text{LLM},t}(s_{t},a_{t})$
13:    else
14:     $r_{t}\leftarrow r_{\text{env},t}$
15:    end if
16:    Append $(s_{t},a_{t},r_{t},s_{t+1})$ to $\tau_{\text{episode}}$
17:  end for
18:  Store $\tau_{\text{episode}}$ in $\mathcal{D}$
19:  $R_{\text{episode}}\leftarrow\sum_{t}r_{\text{env},t}$; update window $W$
20:  if $|W|\geq 5$ then $\triangleright$ Difficulty update
21:    $\bar{R}\leftarrow\text{mean}(W)$
22:    Update $d$ using thresholds $1.05\tau$ and $0.95\tau$ with hysteresis and anti-stagnation
23:  end if
24:  if A7 and regeneration trigger fired then
25:    $(r_{\text{LLM}},\pi_{\text{LLM}})\leftarrow\text{LLM-Generate}(d,W,\text{task})$
26:    Validate templates via syntax, sandbox, and safety checks
27:  end if
28:  if $|\mathcal{D}|\geq B_{\text{batch}}$ then $\triangleright$ Inner loop
29:    for $k=1$ to $K_{\text{updates}}$ do
30:     Sample minibatch $\mathcal{B}$; update critic by TD error
31:     Compute actor loss $\mathbb{E}[Q_{\phi}(s,\pi_{\theta}(s))]$
32:     if A7 then
33:      Add MSE regularization $\alpha(t)\|\pi_{\theta}(s)-\pi_{\text{LLM}}(s)\|_{2}^{2}$
34:     end if
35:     Update $\theta,\phi$ and their targets via soft updates
36:    end for
37:  end if
38:end while
4.4 LLM-Generated Reward Shaping (A7)
A7 extends A2 by adding LLM-shaped reward and prior policy regularization.
Template-based reward generation.
At regeneration events (difficulty change $\Delta d\geq 0.15$, 10 episodes since last generation, or template switch, with a minimum of 3 episodes between generations), GPT-4o-mini is prompted with the current difficulty, performance statistics, and high-level goals. Instead of emitting raw Python code, the LLM outputs parameters for pre-validated templates (e.g., weights over delay, queue length, phase switching). The templates enforce weight normalization, bound outputs, and prevent NaNs/Infs. A three-stage validation pipeline (syntax, sandboxed execution, safety checks) filters out invalid generations. If three consecutive generations fail, the system falls back to environment-only rewards until the next trigger.
Prior policy regularization.
The LLM also parameterizes a prior policy $\pi_{\text{LLM}}(a|s)$, represented as logits over actions. This prior is never used for action selection or initialization; the actor $\pi_{\theta}$ always chooses actions and starts from random weights. The prior is first generated after an exploration episode and then used to regularize the actor via:
$\mathcal{L}_{\text{reg}}(s)=\alpha(t)\,\|\pi_{\theta}(s)-\pi_{\text{LLM}}(s)\|_{2}^{2},$
with $\alpha(t)$ linearly decaying from $0.5$ at episode 1 to 0 at episode 200. The full actor loss is:
$\mathcal{L}_{\text{actor}}=-\mathbb{E}_{s\sim\mathcal{B}}[Q_{\phi}(s,\pi_{\theta}(s))]+\mathbb{E}_{s\sim\mathcal{B}}[\mathcal{L}_{\text{reg}}(s)].$
This encourages early exploration along semantically meaningful behaviors (e.g., maintaining green waves, prioritizing high-pressure links) while gradually handing over control to the learned policy.
Difficulty-dependent weighting.
The auxiliary reward weight is:
$w(d)=0.1+0.4d,$
where $d$ is a normalized difficulty score from a dedicated analyzer. Given the observed range $d\in[0.20,0.69]$, $w(d)\in[0.18,0.38]$, ensuring modest influence at easy levels and stronger shaping in harder regimes.
The combined reward is:
$r_{t}=\begin{cases}r_{\mathrm{env},t},&\text{A8},\\[3.0pt] (1-w(d))\,r_{\mathrm{env},t}+w(d)\,r_{\mathrm{LLM},t},&\text{A2, A7}.\end{cases}$
5 Experimental Design and Evaluation
We assess MAESTRO by (i) isolating the contribution of LLM-shaped rewards relative to an adaptive curriculum with template rewards, and (ii) evaluating stability and efficiency in a non-stationary TSC environment.
Conditions.
We train:
  • A2 (Curriculum Baseline): llm_adaptive curriculum, template reward.
  • A8 (Alternative Curriculum): llm curriculum, template reward.
  • A7 (Full MAESTRO): llm_adaptive curriculum, LLM-shaped rewards and prior regularization.
All use identical MADDPG architectures and hyperparameters.
Comparisons.
A2 vs. A7 isolates reward shaping; A2 vs. A8 compares curriculum modes. Each configuration is trained for 200 episodes with four seeds.
5.1 Environment and Curriculum Settings
We use the HZ network with 16 intersections; each episode has 360 steps. We report episode return and standard TSC metrics. The difficulty update mechanism is as in Section 4.2, with difficulty clipped to $[0.3,1.0]$. Difficulty trajectories differ by seed due to performance-based updates.
In A7, difficulty changes and stagnation events also trigger LLM reward regeneration; A2 and A8 share the same curriculum logic but never generate rewards.
5.2 Metrics
We evaluate:
  • Traffic efficiency: mean delay, queue length, throughput, and cumulative return (episodes 180–199).
  • Learning dynamics: episodes to reach target performance (165.0), time to reach 90% of final return, variance and coefficient of variation (CV) of late-training returns, and a Sharpe-like risk-adjusted metric.
  • Curriculum behavior: difficulty trajectories, counts of adjustments, and final difficulty levels.
5.3 Protocol
We run each configuration with seeds 200, 300, 400, and 500. Training consists of 200 episodes (72k steps). Every second episode we run a deterministic evaluation. Random seeds control PyTorch, NumPy, Python RNG, and CuDNN. In A7, LLM template acceptance is deterministic given the seed.
For each configuration and seed, we summarize final performance by the mean return over episodes 180–199. Pairwise comparisons use Mann–Whitney U tests with Bonferroni correction; effect sizes use Cohen’s $d$, and uncertainties use 95% bootstrap confidence intervals over seeds.
5.4 Visualization
We report four plots: (1) learning curves with 95% confidence bands (Fig. 3); (2) a radar chart over normalized metrics (Fig. 4); (3) ablation bar plots for mean and variance (Fig. 5); and (4) a CV comparison (Fig. 6).
6 Results
Figure 3 shows training curves; Table 3 summarizes late-training returns.
6.1 Traffic Efficiency
Both LLM-enhanced configurations (A7, A8) improve traffic metrics over A2. Table 1 reports seed 200 as a representative example. A8 yields the largest improvements, reducing mean delay by 27.8% and increasing throughput by 9.2%; A7 also significantly reduces delay (19.9%) and wait time.
Table 1: Late-training traffic metrics (seed 200). Higher throughput and lower times are better. Percentages are relative to A2.
Condition Throughput Avg Travel Avg Delay Avg Wait
(veh/ep) Time (s) (s) Time (s)
A2 (LLM curr. + template) 1973 793 472 502
A7 (Full MAESTRO) 2090 (+5.9%) 739 ($6.9\%$) 378 ($19.9\%$) 422 ($15.8\%$)
A8 (Alt. curr. + template) 2155 (+9.2%) 708 ($10.8\%$) 341 ($27.8\%$) 383 ($23.6\%$)
Table 2: Traffic efficiency metrics averaged across all four seeds (200, 300, 400, 500) during late-training episodes (180–199). Values reported as mean $\pm$ standard deviation across seeds. Higher throughput and lower temporal metrics indicate better performance. Percentages denote relative changes compared to A2 baseline.
Condition Throughput Avg Travel Avg Delay Avg Wait
(veh/ep) Time (s) (s) Time (s)
A2 (LLM Curr. + Template) $1570\pm 212$ $1030\pm 131$ $711\pm 143$ $763\pm 127$
A7 (Full MAESTRO) $1776\pm 188$ (+13.1%) $919\pm 116$ ($10.8\%$) $571\pm 141$ ($19.7\%$) $640\pm 127$ ($16.2\%$)
A8 (Alt. Curr. + Template) $1745\pm 302$ (+11.1%) $954\pm 169$ ($7.4\%$) $602\pm 203$ ($15.4\%$) $652\pm 215$ ($14.5\%$)
Figure 3: Training performance over 200 episodes. Lines show mean episode return; bands show 95% confidence intervals. A7 exhibits the narrowest band, indicating high stability across seeds.
6.2 Final Performance and Stability
We use episode return (cumulative reward) averaged over episodes 180–199 as the primary metric. Table 3 presents summary statistics.
Table 3: Late-training episode returns (higher is better). CV: coefficient of variation. Sharpe Ratio (SR) = (mean – A2 mean)/std.
Condition Curric. Reward Mean $\pm$ SD CV (%) SR Best Worst
A2 (Baseline) llm_adaptive Template $157.04\pm 3.07$ 1.95 159.63 152.82
A7 (Full MAESTRO) llm_adaptive LLM-shaped $158.87\pm 1.20$ 0.76 1.53 160.15 157.76
A8 (Alt. curric.) llm Template $160.23\pm 4.73$ 2.95 0.67 166.08 154.72
A8 attains the highest mean and best seed but also the highest variance. A7 achieves moderate performance gains over A2 with dramatically lower variance. Its worst seed exceeds the A2 mean, making it strongly preferable when reliability matters. Mann–Whitney U tests are underpowered at $n=4$ seeds but effect sizes are large; we treat these results as practical rather than formal significance.
Figure 4: Normalized comparison across five metrics (0–1 scale). A7 balances final return, stability, learning speed, sample efficiency, and robustness. A8 dominates in peak return but is less stable; A2 is consistently weaker.
6.3 Ablation Insights
Curriculum mode (A2 vs. A8).
Switching from llm_adaptive to llm increases mean return but also variance. More expressive curricula clearly help exploration but can destabilize training.
Reward shaping (A2 vs. A7).
Introducing LLM-shaped reward and prior regularization under the same curriculum yields consistent gains and a large reduction in variance. This supports the view that reward templates act as structured priors, smoothing exploration and updates.
Stability–performance trade-off (A7 vs. A8).
A7 and A8 trace a classic trade-off: A8 is best for peak performance, A7 for risk-adjusted performance. For real deployments, A7’s reliability and lower CV are more attractive.
Figure 5: Mean returns and standard deviations. A8 yields the largest raw gain; A7 offers smaller gains with much tighter error bars.
[…]
7 Discussion and Future Work
7.1 Key Findings
Generative shaping improves stability.
[…]
Curriculum is the main performance lever.
The largest jump in mean return arises from changing the curriculum mode (A2 $\rightarrow$ A8), underscoring that task selection and sequencing dominate asymptotic performance. LLM-generated semantically rich contexts make curricula a high-impact design dimension.
Small return gains, large domain impact.
While episode return improvements appear modest (1–2%), they correspond to substantial domain gains: 20–28% reductions in delay and 6–9% increases in throughput. This highlights the importance of reporting task-level metrics, not just abstract returns.
7.2 Limitations and Directions
Domain generality.
Our evaluation is limited to the HZ TSC network. Extending MAESTRO to other cooperative MARL domains (multi-robot coordination, resource allocation, network routing) is necessary to establish generality.
Model and prompt dependence.
We use GPT-4o-mini with a single prompt design. Future work should explore prompt tuning, alternative LLMs, and few-shot examples to further improve curriculum and reward quality.
Richer training signals.
Currently, MAESTRO shapes rewards and regularizes policies; future extensions could include LLM-informed value initialization, hierarchical options, or dynamic architecture adaptation.
Reward safety and interpretability.
Although template constraints and validation provide strong safeguards, subtle misalignment remains possible. For deployment, human-in-the-loop review, formal verification, or runtime monitoring will be important.
8 Conclusion
We introduced MAESTRO (Multi-Agent Environment Shaping through Task and Reward Optimization), a contextual LLM-aided MARL framework that uses an LLM as an offline training architect rather than an online controller. By combining an adaptive semantic curriculum with template-based LLM reward shaping and prior policy regularization on top of MADDPG, MAESTRO improves traffic efficiency on a realistic 16-intersection network while preserving low-latency deployment.
A controlled ablation across three LLM-assisted configurations yields two core insights: (i) the LLM’s curriculum generation is the dominant lever for performance, and (ii) constrained, template-based reward shaping significantly improves stability and risk-adjusted returns. Modest gains in episode return translate into large practical improvements in TSC metrics, including up to 28% delay reduction and 9% throughput increase. Our results suggest that LLMs are particularly well-suited as architects of training processes, where their semantic reasoning can structure curricula and rewards for efficient, robust MARL, while leaving execution to domain-appropriate learned controllers.
Appendix A Detailed Experimental Results
This appendix provides comprehensive experimental details and supplementary analyses referenced in Section 6.
A.1 Per-Seed Performance Breakdown
Table 4 provides complete per-seed performance data over episodes 180–199. A8 exhibits the widest performance range (11.36 points), with seed 400 achieving the best result (166.08) but seed 500 falling below baseline (154.72). In contrast, A7 shows strong consistency across all four seeds (range: 2.39 points), with all seeds exceeding baseline.
Table 4: Per-seed traffic control performance (episode return over episodes 180–199). Training rewards extracted from simulation logs. All configurations completed 4 seeds each.
Condition Seed 200 Seed 300 Seed 400 Seed 500 Mean SD Range
A2 (Baseline) 158.97 156.75 159.63 152.82 157.04 3.07 6.81
A7 (LLM Rewards) 159.63 157.93 160.15 157.76 158.87 1.20 2.39
A8 (Alt Curriculum) 161.18 158.95 166.08 154.72 160.23 4.73 11.36
Per-Seed Range 2.66 2.20 5.93 4.94
Risk Analysis.
A7’s worst-case performance (157.76) exceeds the baseline mean (157.04) by 0.72 points, while A8’s worst-case (154.72) falls 2.32 points below baseline. This represents a critical difference: A7 guarantees improvement regardless of random seed, whereas A8 carries 25% risk of degradation.
Seed-Level Deviations.
Computing per-seed deviations from respective configuration means:
  • A2: $\pm$[1.93, 0.29, 2.59, 4.22] points (max deviation: 4.22)
  • A7: $\pm$[0.76, 0.94, 1.28, 1.11] points (max deviation: 1.28)
  • A8: $\pm$[0.95, 1.28, 5.85, 5.51] points (max deviation: 5.85)
[…]
A.2 Curriculum Adaptation Dynamics
Figure 7: Curriculum difficulty trajectories over 200 episodes for representative seeds. Both A2 and A7 (sharing llm_adaptive mode) exhibit bidirectional adaptation: difficulty increases when performance improves but decreases after performance drops. Shaded regions indicate episodes where difficulty was reduced to prevent overchallenge. Note: A8 uses llm mode with different adaptation characteristics and is not plotted here.
Figure 7 plots curriculum difficulty scores against training episodes for seeds 200. Key observations:
A2 and A7 Adaptation Patterns.
Both configurations using llm_adaptive curriculum show similar adaptation dynamics:
  • Bidirectional adjustments: Difficulty increases by $\sim$0.05–0.10 after performance gains but decreases by similar magnitude after performance drops.
  • Adaptation frequency: Curriculum updates occur approximately every 2–3 episodes, with $\sim$80 total updates over 200 episodes.
  • Difficulty range: Difficulty scores span [0.3, 0.7], indicating moderate challenge levels throughout training.
A8 Curriculum Behavior.
The llm curriculum mode in A8 exhibits different characteristics, though detailed trajectory analysis is needed for complete characterization. Preliminary observations suggest less frequent but larger difficulty adjustments compared to llm_adaptive.
Performance-Difficulty Correlation.
Overlaying episode returns with difficulty trajectories reveals that temporary difficulty reductions often precede renewed performance gains (typically 5–10 episodes after reduction). This supports the hypothesis that adaptive curricula prevent premature convergence by providing temporary "easier" scenarios when agents struggle.
A.3 Statistical Test Details
Table 5: Pairwise statistical comparisons using Mann–Whitney U tests on seed-level episode returns. All tests use $\alpha=0.05$ significance level.
Comparison U Statistic p-value Cohen’s d Effect Size Significant?
A2 vs A7 1.0 0.476 0.81 Large No
A2 vs A8 2.0 0.343 0.89 Large No
A7 vs A8 0.5 0.381 0.43 Medium No
Interpretation.
While effect sizes (Cohen’s $d>0.8$ for A2 vs A7/A8) indicate practically meaningful differences, limited sample size ($n=4$ seeds) prevents statistical significance at $\alpha=0.05$. Power analysis reveals that detecting these effect sizes with 80% power requires:
  • A2 vs A7: $n\geq 12$ seeds per condition
  • A2 vs A8: $n\geq 10$ seeds per condition
  • A7 vs A8: $n\geq 20$ seeds per condition (lower effect size)
Practical Significance.
[…]
A.4 Computational Cost Breakdown
Table 6: Detailed computational cost analysis for 200-episode training run on single NVIDIA RTX-class GPU.
Component Time per Call Call Frequency Total Time Percentage
CityFlow Simulation 2.8 hours 80.0%
MADDPG Training 0.65 hours 18.6%
LLM Curriculum Calls 2.3s Every 2–3 eps 2.5 minutes 1.2%
LLM Reward Calls (A7) 1.8s Every 5–10 eps 1.0 minutes 0.5%
Total 3.5 hours 100%
API Cost Details.
Using GPT-4o-mini pricing (as of experiment date):
  • Curriculum generation: $\sim$80 calls $\times$ \$0.004/call = \$0.32
  • Reward generation (A7): $\sim$40 calls $\times$ \$0.004/call = \$0.16
  • Total per run: \$0.47 (A7), \$0.32 (A2/A8)
For a full 4-seed ablation study: $\sim$\$1.88 (A7 with 4 seeds) + \$1.28 (A2 with 4 seeds) + \$1.28 (A8 with 4 seeds) = \$4.44 total.
Economic Feasibility.
At scale, the \$0.47 incremental cost for LLM-enhanced training is negligible compared to deployment benefits. For a city-wide traffic system, even 1% improvement in throughput or delay reduction translates to significant economic value (reduced fuel consumption, lower emissions, improved quality of life).
A.5 Traffic Metrics: Multi-Seed Analysis
Table 7: Traffic efficiency metrics extracted from simulation logs (late-training episodes, seed 200 only). Single-seed results shown; future work will aggregate across all seeds for stronger statistical support.
Condition Throughput Avg Travel Time Avg Delay Avg Wait Time
(vehicles/ep) (seconds) (seconds) (seconds)
A2 (Baseline) 1973 793 472 502
A7 (LLM Rewards) 2090 (+5.9%) 739 ($6.9\%$) 378 ($19.9\%$) 422 ($15.8\%$)
A8 (Alt Curriculum) 2155 (+9.2%) 708 ($10.8\%$) 341 ($27.8\%$) 383 ($23.6\%$)
Real-World Impact.
For a 6-minute (360 timestep) episode:
  • A7 delay reduction: 94 seconds per vehicle (472 $\rightarrow$ 378s)
  • A8 delay reduction: 131 seconds per vehicle (472 $\rightarrow$ 341s)
  • Throughput increase: 117–182 additional vehicles per episode
Scaling to 16 intersections over 1 hour (10 episodes):
  • A7: $\sim$1,170 additional vehicle trips, saving 15.6 vehicle-hours of delay
  • A8: $\sim$1,820 additional vehicle trips, saving 21.8 vehicle-hours of delay
Note on Multi-Seed Data.
Current analysis uses seed 200 only due to time constraints. Future work will extract and aggregate traffic metrics across all seeds (seeds 200, 300, 400, 500 for all configurations A2, A7, A8) to provide mean $\pm$ SD for each metric, strengthening statistical rigor.
A.6 Learning Curve Analysis
Figure 8: Learning curves for A2, A7, and A8 configurations over 200 episodes. Solid lines show mean episode return; shaded regions indicate standard deviation across seeds. A7 exhibits tightest confidence bands, confirming low variability.
Convergence Characteristics.
  • A2 (Baseline): Converges to $\sim$157 by episode 120, stable thereafter
  • A7 (LLM Rewards): Converges to $\sim$159 by episode 110 (10 episodes earlier), maintains narrow confidence bands (SD << 2 points)
  • A8 (Alt Curriculum): Highest variance throughout training, achieves peak $\sim$160 by episode 140 but shows continued fluctuation
Sample Efficiency.
A7 reaches baseline performance (156.93) approximately 15 episodes earlier than A2, suggesting improved sample efficiency from LLM reward shaping. However, this advantage is modest and requires further investigation with more seeds.
A.7 Additional Risk Metrics
Table 8: Comprehensive risk-adjusted performance metrics. Sharpe Ratio = (mean - baseline) / std; Sortino Ratio uses downside deviation; VaR = Value at Risk (5th percentile).
Condition Sharpe Ratio Sortino Ratio VaR (5%) Max Drawdown Success Rate
A2 (Baseline) 152.82 4.22 75% (3/4)
A7 (LLM Rewards) 1.53 2.16 157.76 1.28 100% (4/4)
A8 (Alt Curriculum) 0.67 1.12 154.72 5.51 75% (3/4)
Interpretation.
[…]
  • Sortino Ratio: A7’s 2.16 vs A8’s 1.12 shows even larger advantage when considering only downside risk
  • VaR (5%): A7’s worst case (157.76) is the highest among all configurations, indicating lowest tail risk
[…]
  • Success Rate: A7’s 100% vs A8’s 75% confirms zero deployment risk
These comprehensive risk metrics consistently support A7 as the optimal choice for production deployment where reliability is paramount.
A.8 Configuration Details
For complete reproducibility, we document exact experimental configurations:
A2 (Baseline).
  • Curriculum: LLM-generated, mode = llm_adaptive
  • Reward: Template-based (flow + wait + pressure)
  • Seeds: 200, 300, 400, 500 (all completed)
  • Log files: logs/maddpg_baseline_200ep_adaptive_*
A7 (LLM Rewards).
  • Curriculum: LLM-generated, mode = llm_adaptive (same as A2)
  • Reward: LLM-generated with difficulty-adaptive weighting, $w(d)=0.1+0.4\cdot d$ for normalized difficulty $d\in[0,1]$
  • Seeds: 200, 300, 400, 500 (all completed)
  • Log files: logs/llm_maddpg_200ep_enhanced_V2_adaptive_*
  • Regenerations: $\sim$80 reward updates over 200 episodes
A8 (Alternative Curriculum).
  • Curriculum: LLM-generated, mode = llm (different from A2/A7!)
  • Reward: Template-based (same as A2)
  • Seeds: 200, 300, 400, 500 (all completed)
  • Log files: logs/llm_enhanced_FIXED_seed*
All experiments use MADDPG with consistent hyperparameters (learning rate, batch size, etc.) and 360-timestep episodes on Hangzhou 16-intersection network.